40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 14 13:29:51 2018, gautam, Update, General, PSL shutter closed, all optics misaligned 
    Reply  Thu Jun 14 19:30:12 2018, gautam, Update, General, IFO alignment restored 
       Reply  Thu Jun 14 22:45:05 2018, johannes, Update, General, AUX beam SRC alignment 
          Reply  Sat Jun 16 00:26:48 2018, gautam, Update, General, PRC modescan attempt 
             Reply  Sat Jun 16 01:25:29 2018, Koji, Update, General, PRC modescan attempt 
                Reply  Sat Jun 16 20:57:59 2018, Jon, Update, General, PRC modescan attempt 40M_SRMI_AM_annotated.pdf40M_DRMI_AM.pdf
Message ID: 13967     Entry time: Thu Jun 14 19:30:12 2018     In reply to: 13962     Reply to this: 13968
Author: gautam 
Type: Update 
Category: General 
Subject: IFO alignment restored 

All optics have been re-aligned. Jon/Johannes will elog about the work today.

ELOG V3.1.3-